OpenStudy (rayman5766):

What is the ratio of x to y? 3/5 = x/x+2y A-3 B -1 /3 C 1/3 D 3

OpenStudy (amistre64):

finding x/y would seem to account for it

OpenStudy (amistre64):

mutliply thru by 5(x+2y)

OpenStudy (rayman5766):

I was kinda thinking B

OpenStudy (amistre64):

3(x+2y) = 5x should be simple to work out a solution to x/y

OpenStudy (rayman5766):

ok first things first we have to find out what is x and y

OpenStudy (amistre64):

we dont have to find out what x or y is, we simply have to solve the setup for x/y on one side

OpenStudy (rayman5766):

oh k that makes since

OpenStudy (amistre64):

3(x+2y) = 5x 3x + 6y = 5x 6y = 2x

OpenStudy (amistre64):

what is x/y? divide by 2 and divide by y